# Network Analytics for Anti-Money Laundering – A Systematic Literature Review and Experimental Evaluation

### v2

@@ Abstract
Money laundering presents a pervasive challenge, burdening society by financing illegal activities. To more effectively combat and detect money laundering, the use of network information is increasingly being explored, exploiting that money laundering necessarily involves interconnected parties. This has lead to a surge in literature on network analytics (NA) for anti-money laundering (AML). The literature, however, is fragmented and a comprehensive overview of existing work is missing. This results in limited understanding of the methods that may be applied and their comparative detection power. Therefore, this paper presents an extensive and systematic review of the literature. We identify and analyse 97 papers in the Web of Science and Scopus databases, resulting in a taxonomy of approaches following the fraud analytics framework of Bockel-Rickermann et al. [17]. Moreover, this paper presents a comprehensive experimental framework to evaluate and compare the performance of prominent NA methods in a uniform setup. The framework is applied on the publicly available Elliptic data set and implements manual feature engineering, random walk-based methods, and deep learning GNNs. We conclude from the results that network analytics increases the predictive power of the AML model with graph neural networks giving the best results. An open source implementation of the experimental framework is provided to facilitate researchers and practitioners to extend upon these results and experiment on proprietary data. As such, we aim to promote a standardized approach towards the analysis and evaluation of network analytics for AML.
@@ 1 Introduction
Money laundering is the process of concealing illegally obtained funds by passing them through a complex sequence of transactions. Laundered money can then be used in the legitimate financial system to fund and facilitate criminal activities, such as drug trafficking, corruption, and terrorism [70]. Yearly, around $2\%$ to $5\%$ of global GDP is laundered worldwide, amounting to USD 2 trillion, as estimated by the United Nations Office on Drugs and Crime (UNODC) [116]. This illegally obtained money needs to enter the financial system at some point [70]. Therefore, legislative bodies have put stringent rules on financial institutions, requiring them to have safeguards in place and to report suspicious transactions/activities to the Financial Intelligence Unit (FIU). Despite record fines and mounting pressure by regulators, money laundering represents a growing problem [123]. Anti-money laundering practices are largely ineffective [33], with Europol estimating only 1.1% of criminal profits in the EU being confiscated [35].
@@
Recently, there has been increasing attention to leverage network information for anti-money laundering (AML) efforts, seeking to exploit the fact that money laundering typically involves transactions between interconnected parties. Money laundering generally requires three main steps [116, 70]. The first is placement, where dirty money is integrated into the financial system. The second is layering, where money is transferred and mixed to obscure its illegal origin. The third is integration, where the laundered money is integrated into wider circulation by doing seemingly legitimate purchases. Since these steps all involve multiple parties, the adoption of network analytics holds promising potential [12].
@@
The increasing attention has resulted in a growing body of literature covering network analytics for anti-money laundering. However, the swift expansion of the field has simultaneously resulted in a lack of a comprehensive overview. Most literature reviews on financial fraud barely discuss anti-money laundering using network analytics [91, 69, 39, 53]. Therefore, as a first step, this paper aims to supplement the literature by providing a systematic literature review of network analytics (NA) for anti-money laundering. To this end, we constructed queries in an iterative way to identify relevant papers from Web of Science (WOS) and Scopus, resulting in 97 papers.
@@
The papers were analysed according to different dimensions. First, we looked at the evolution of the number of published papers over time. Second, each paper was classified according to five categories; publication data, the applied learning method, the evaluation metrics used, the main objective, and the kind of data used. We identify an increasing focus on crypto currencies. Our work compares the differences between those papers and the research covering classic financial transactions. Third, the top-cited work is reviewed via the framework introduced in [17].
@@
Our literature review and subsequent analysis reveals that multiple supervised, unsupervised and visualisation methods have been developed and applied. Nevertheless, due to the literature being fragmented and these methods mostly addressing only a part of the AML problem, there is limited insight into how different methods compare and which work best. Furthermore, most state-of-the-art network analytics methods have not yet found main-stream adoption for anti-money laundering. Therefore, this work provides a comprehensive experiment, evaluating the most prominent methods.
@@
  • We provide an extensive and critical systematic literature review on network analytics for anti-money laundering, to gain both comprehensive and deep insight;
  • We apply additional network analytics methods to anti-money laundering;
  • We conduct a comprehensive experimental evaluation of the key methods found in the literature on the open-source Elliptic [1, 122] data set;
  • We implement methods in a uniform manner and facilitate replication of the presented results by providing public access to the code at https://github.com/B-Deprez/AML_Network, aiming to promote a standardized approach towards analysing the use of NA for AML and to encourage further research on network analytics for anti-money laundering.
This paper is structured as follows. We present the methodology for the structured literature review in Section 2, whereas Section 3 presents the results of the review. Section 4 describes the set-up of the empirical evaluation, whereas the results are presented in Section 5. Finally, overall conclusions are presented in Section 6.
@@ 2 Literature Review
In this section, we present our methodology for conducting a comprehensive review of the literature on NA for AML. First, we establish the scope of the review, as well as the queries that are used and the processing of retrieved papers. Then, we provide the results.
@@ 2.1 Methodology
We iteratively defined a broad search using different queries for finding the relevant papers. We considered both network and graph as two near-synonyms used by researchers [8, 16, 47]. Additionally, we included papers using analytics and/or analysis. Finally, the term anti-money laundering is not broad enough, since methods might try to, e.g., “detect money launderers”. Therefore, we used the following queries:
  • graph analy*ANDmoney launder*
  • network analy*ANDmoney launder*
These queries are used to search for papers on WoS and Scopus, since these databases have a very high combined degree of completeness, and both are well respected for usage in research [88, 17]. This search is limited to papers published before 2023. Publishing year 2023 and later are, at the time of writing, too recent to judge the acceptance and appreciation by the research community, and hence the paper’s overall quality.
We first perform a title and abstract scan to decide upon the final set of papers. Second, papers are filtered using a full text scan. We exclude, among others, those dealing with law, or those not applying network analytics to anti-money laundering. A few papers were retrieved that only briefly mentioned money laundering in their abstract/introduction, but did not deal with AML as part of the research.
Once the final selection of relevant papers is obtained, we carefully analyse and classify them according to five categories; publication data, the applied machine learning methods, the evaluation metrics, the main objective, and the kind of data. These categories have sub-categories to obtain meaningful classifications; in Table 1. This classification is used to have a structured analysis and capture the main trends when presenting the results in Section 2.2 and discussing the literature in Section 3.
@@
Learning Method Supervised Application of supervised method
@@
Table 1: Paper categorisation. The (sub-)categories with explanation for the classification of the literature.
This work provides both a comprehensive and in-depth review of the literature. The comprehensive review discusses the general trends, based on the categories in Table 1, for all papers in scope. The in-depth review discusses the top-cited paper in more detail to get insights into the current state-of-the-art. This state-of-the-art contains the top-$10\%$ cited per year, and the top-$10\%$ cited overall, according to Google Scholar. The absolute number of papers are determined by rounding up the obtained number of papers; $\lceil 0.10\cdot n\rceil,$ with $n$ the total number of papers. This provides a view on the research that has been well received.
As part of the in-depth analysis, we discuss the scope of all review papers selected. These give an additional overarching view of the literature. Additionally, this allows us to identify gaps in those reviews, showing how our work complements and extends the existing literature.
@@
Figure 1: Review overview. The search and filtering method with the number of remaining papers between brackets.
@@ 2.2 Results
The initial number of papers for all queries is summarised in Figure 1. The first numbers are presented for WoS and Scopus separately, meaning that duplicates are present. These papers were filtered, as described in the previous section, and duplicate papers were removed, resulting in 97 papers, 11 of which are review papers. Table LABEL:tab:literature gives the overview, using the classification from Table 1. Since the scope of the review papers is broader than for individual papers, no category-specific information is given.
Figure 2: Research Output. The number of papers incorporated in the full query and after filtering per year.
Figure 2 shows the evolution of the number of papers before and after filtering. An increasing trend in research output is observed. Similar figures are provided in B where the results are presented per category from Table 1. These category-specific figures are compiled without the review papers. Next to these plots, Table 2 summarises the journals and conference proceedings, including only the venues having more than one publication.
We see in Table LABEL:tab:literature that 26 papers specifically deal with crypto currencies. Therefore, C represents crypto-related research separately, since the setting of AML is different from normal financial transactions. This is also visible in the presented plots.
Finally, D contains the results for a top percentage of papers. The resulting 20 papers are presented in Table 3, where we indicate if the paper is a review and/or crypto-related paper. The distribution among the categories is in line with the overall literature. Additionally, $25\%$ of these papers are on crypto currency, which is similar to the $26.5\%$ seen in the full scope.
@@
Table 2: Venues. Journals/conferences that have two or more publications in scope of this review.
@@ 3 Discussion on the Literature
3.1 Full Scope of Papers
Figure 2 shows an increasing interest in the literature in applying network analytics for anti-money laundering, especially after 2010, with an average increase in output of $61\%$ year-on-year for the papers in scope.
The methods proposed in the papers are supervised, unsupervised, semi-supervised and visualisations. The most popular method is unsupervised learning, followed by supervised learning and visualisations, as shown in Figure 1. This may be due to money laundering being very uncommon, resulting in highly skewed data sets with very few labelled observations. In addition, criminals try to evade detection by changing tactics. The performance of (supervised) models trained on historical data can therefore decrease over time [17].
Additionally, we remark that most supervised methods could be classified as semi-supervised. Since not all money laundering transactions are detected, there are cases in the legitimate data that are actually money laundering. However, we follow the classification as represented in the papers, not using our personal interpretation of the problem.
We identified multiple objectives of the methods for anti-money laundering. Figure 2 shows that the objective is most often detecting suspicious clients, followed by detecting suspicious transactions, money flows and communities. Client classification combines payments made over a period of time and the person’s characteristics. This seems better for detecting money laundering, since it is often done over a longer period, with the aim to make the individual transaction appear normal. Therefore, models trained on a series of transactions, leveraging an individual’s behaviour, might obtain higher performance. Figure 2 shows that methods for crypto-currency are most often introduced for suspicious transaction detection. Due to the pseudo-anonymity, it is much harder to know which wallets belong to one person [140, 133], making it less feasible to build profiles.
In recent years, more studies have been done on flow/chain detection, tracking payments over multiple steps/people. Although this requires more computing power, it more closely resembles reality, since money laundering is done over multiple steps [116, 70]. Hence, methods leveraging these different steps and actors are becoming more widely used.
The most popular metrics to evaluate the methods are the recall, $F1$-score, precision and time, as shown in Figure 3. We see that training/prediction time was a popular method in earlier studies, but relatively fewer papers use time in more recent work, which might be a consequence of the increase in computing power. Additionally, the popularity of the precision, recall and $F1$-score is a recent and increasing trend in the literature.
We remark two surprising results in our analysis. First, only one paper used the area under the precision-recall curve (AUC-PR). Due to the high label imbalanced, this metric is shown to be more appropriate than the area under the ROC curve (AUC-ROC) for evaluating performance [3, 94]. Therefore, we would expect it to be more widely used. Second, some papers solely reported accuracy [29, 129, 45, 131]. Given the class imbalance and the dependence on a threshold, accuracy however fails in providing an appropriate assessment of performance.
Finally, Figure 4 shows the data used. Most data sets are open-source, although proprietary data is also used in almost $50\%$ of the papers. Only a minority of papers report their findings using synthetic data. There is, however, a major difference between crypto- and other research. Looking at Figure 1(a), we see that crypto-research is almost exclusively done via open-source data. This results from the blockchain’s freely accessible distributed ledger [122]. On the other hand, classic transaction data cannot be shared, due to privacy reasons, resulting in half the data sets being proprietary.
Following this last remark, one would expect to have (almost) no research done on open-source data sets, for non-crypto-payments. Figure 1(a) shows that almost $30\%$ of non-crypto related papers use open-source data. These papers can be categorised in four groups:
  • Methods that are evaluated on non-financial network data, whether or not next to proprietary/synthetic money laundering data: [139, 93, 84, 99]
  • Methods evaluated on government data, assumed to be open-source: [27, 58, 119, 13, 79]
  • Methods constructed to analyse leaked documents/scandals, e.g., the Panama papers: [63, 64, 76, 4, 124, 55, 24]
  • Other: [89, 15, 19]
For the Other category, we have the following. The data in [89] is said to be publicly available, without giving any more information. The data in [15] is scraped from Hack Forums. The data in [19] is a Kaggle data set that is based on synthetic data. We categorise this data set open-source, since it is made publicly available.
@@ 3.2 Review Papers
The research in [91, 68, 75] focuses on financial fraud in general, resulting in less emphasis on anti-money laundering or graph learning. The authors of [91] construct a framework to classify financial fraud, but only one paper on money laundering was included, which was also the only paper on network analytics. The authors in [68] give a high-level overview of the general difficulties of machine learning solutions for fighting financial crime. They noted that network-based methods have a lot of potential to process many transactions handled by financial institutions. The paper is closer to a discussion paper than to a systematic literature review. The scope of the review in [75] is limited. It deals with both credit card fraud and money laundering, but only incorporates visualisation methods used for outlier detection.
@@
Other review papers deal specifically with AML [69, 104, 39]. The work in [69] is most similar to ours. The authors give an overview of the methods applied in the literature, but their main topic is interpretability. Hence, their focus is less on the overall AML literature. The review in [104] discusses the different methods applied to AML. It touches upon different network methods, but it is not a systematic literature review. The framework constructed in [39] again deals with a wider array of methods, with the authors concluding that network analytics is the most promising. However, the number of papers on network analytics for AML in [39] is limited.
@@
The work in [44] looks at the effects of money laundering policies, by measuring the change in network features when a major AML law is introduced. It analyses cluster size and the degree, closeness and betweenness centrality. Hence, the scope of [44] is different, since we focus on enhancing anti-money laundering using network features.
@@
A significant number of review papers deal with crypto currency related money laundering [59, 26, 6, 53]. The work in [59] describes challenges in detecting money laundering on the blockchain. As a result, the scope is smaller, with a total of 45 unique references in the paper. These different challenges are partially addressed in [6, 53]. The experiment in [6] compares different supervised machine learning methods, with the statistics of the ego-network of the different nodes in the Bitcoin transaction network (Elliptic data set [1, 122]) as features.
@@
Other papers cover a specific money laundering method and propose tailor-made solutions. The review in [53] discusses the application of anomaly detection, based on crime-specific transaction patterns. The paper also covers Ponzi schemes and blackmail campaigns. Their scope is more limited, with a total of 33 references in the paper. In the same line, the research in [26] gives an overview of different characteristics of financial crime, and uses these to construct knowledge graphs for building legal cases.
Compared to the literature reviews presented above and to the best of our knowledge, this paper is novel in multiple ways.
  • This is the first systematic literature review on network analytics methods specifically for anti-money laundering;
  • This work will be the first to construct an experimental evaluation comparing multiple different network analytics methods against each other on real-life money laundering data;
  • Our scope is not limited to a particular network method. Our review deals with supervised, unsupervised as well as visualisation methods. It is also not limited in the methods used for featurizing the network. It covers the full spectrum from centrality measures to deep learning and graph neural network;
  • We incorporate papers dealing with money laundering in financial institutions as well as crypto currency.

### v3

@@ Abstract
Money laundering presents a pervasive challenge, burdening society by financing illegal activities. The use of network information is increasingly being explored to more effectively combat money laundering, given it involves connected parties. This led to a surge in research on network analytics (NA) for anti-money laundering (AML). The literature on NA for AML is, however, fragmented and a comprehensive overview of existing work is missing. This results in limited understanding of the methods to apply and their comparative detection power. Therefore, this paper presents an extensive and unique literature review, based on 97 papers from Web of Science and Scopus, resulting in a taxonomy following a recently proposed fraud analytics framework. We conclude that most research relies on expert-based rules and manual features, while deep learning methods have been gaining traction. This paper also presents a comprehensive framework to evaluate and compare the performance of prominent NA methods in a standardized setup. We apply it on two publicly available data sets, comparing manual feature engineering, random walk-based, and deep learning methods. We conclude that (1) network analytics increases the predictive power, but caution is needed when applying GNNs based on the class imbalance and network topology, and that (2) care should be taken with open-source data as this can give overly optimistic results. The open-source implementation facilitates researchers and practitioners to extend upon the results and experiment on proprietary data, promoting a standardized approach for the analysis and evaluation of network analytics for AML.
@@ 1 Introduction
Money laundering is the process of concealing illegally obtained funds by passing them through a complex sequence of transactions. Laundered money can be used in the legitimate financial system to fund and facilitate criminal activities, such as drug trafficking, corruption, and terrorism [77]. Around $2\%$ to $5\%$ of global GDP is laundered worldwide, amounting yearly to USD 2 trillion, as estimated by the United Nations Office on Drugs and Crime (UNODC) [130].
Money laundering differs from other forms of financial fraud, especially credit card and payment fraud, as it is often done over a longer period and only becomes visible when looking at multiple transactions. It is more actively concealed, since sustained laundering is needed to process continuous streams of money. Credit card fraud, on the other hand, is much more about getting as much money out as fast as possible. Hence, the specific characteristics of money laundering merit dedicated research attention.
Since, this illegally obtained money needs to enter the financial system at some point [77], legislative bodies have put stringent rules on financial institutions [15], requiring them to have safeguards and to report suspicious transactions/activities to the Financial Intelligence Unit (FIU). Despite record fines and mounting pressure by regulators, money laundering represents a growing problem [140]. Anti-money laundering practices are largely ineffective [35], with Europol estimating only 1.1% of criminal profits in the EU being confiscated [38].
@@
Money laundering generally requires three main steps [130, 77]. The first is placement, where dirty money is integrated into the financial system. The second is layering, where money is transferred and mixed to obscure its illegal origin. The third is integration, where the laundered money is integrated into wider circulation by doing seemingly legitimate purchases.
Anti-money laundering practices are focused on placement and layering, as after integration via legitimate purchases, the money is in essence successfully laundered. Know-Your-Customer (KYC) protocols screen novel clients when they want to do business with the bank. Afterwards, transactions are continuously monitored to see if a client is involved in money laundering.
Anti-money laundering practices are often a combination of rule-based and risk-based systems [57, 123]. Rule-based systems consist of a collection of different IF-THEN rules. These are defined based on regulatory requirements and expert knowledge [110]. In practice, rule-based systems are still widely used [41, 22, 110]. The main drawback is that many rules are needed as money laundering is complex, which are hard to maintain. Risk-based systems apply more sophisticated methods. Past literature has shown that popular data analytics and machine learning methods have been applied for anti-money laundering [100, 57, 15, 76]. However, Le Khac and Kechadi [76] claim that most commercial tools still rely on more simple statistics to find outliers.
Placement, layering and integration involve multiple parties, and individual transactions often appear normal [15]. Therefore, network analytics is required to uncover the specific money laundering characteristics [15, 10]. The need for network analytics was already mentioned three decades ago by Senator et al. [116], where the authors used visualisations of relations to support experts investing suspicious cases. Although Ngai et al. [100] said that there was still a lack of research at that time, more attention has been given to money laundering and network analytics in the past years.
@@
The increasing attention has resulted in a growing body of literature covering network analytics for anti-money laundering. The swift expansion of the field has simultaneously resulted in a lack of a comprehensive overview. Most literature reviews on financial fraud barely discuss anti-money laundering using network analytics [100, 75, 42, 56]. Therefore, as a first step, this paper aims to supplement the literature by providing a systematic literature review of network analytics (NA) for anti-money laundering. To this end, we constructed queries to identify relevant papers from Web of Science (WOS) and Scopus, resulting in 97 papers.
@@
The papers are analysed according to different dimensions. First, we look at the evolution of the number of published papers over time. Second, each paper is classified according to seven categories; publication data, the method type, the modelling method, the evaluation metrics used, the main objective, the kind of data, and network characteristics. We identify an increasing focus on crypto currencies. Our work compares the differences between those papers and the research covering classic financial transactions. Third, the top-cited work is reviewed via the framework introduced by Bockel-Rickermann et al. [14].
@@
Our literature review and subsequent analysis reveals that multiple supervised, unsupervised and visualisation methods have been developed and applied. Nevertheless, due to the literature being fragmented and these methods mostly addressing only a part of the AML problem, there is limited insight into how different methods compare and which work best. Furthermore, most state-of-the-art (SOTA) network analytics methods have not yet found main-stream adoption for anti-money laundering. Therefore, this work provides a comprehensive experiment, evaluating the most prominent methods.
@@
  • We provide an extensive and critical systematic literature review on network analytics for anti-money laundering, to gain both comprehensive and deep insights, and provide direction for future research.
  • We benchmark a wide range of state-of-the-art network analytics methods in the context of anti-money laundering, comparing the performance of these key methods on two open-source data sets: the Elliptic [37, 139] and the IBM Transactions for AML [6, 36] data set.
  • Based on our findings, we provide insights into the specific challenges of network methods based on the topology of the network and the data generating process.
  • We implement the methods in a uniform manner and facilitate replication of the presented results by providing public access to the code at https://github.com/B-Deprez/AML_Network, aiming to promote a standardized approach towards the analysis and adoption of NA for AML and to encourage further research.
This paper is structured as follows. We present the methodology for the structured literature review in Section 2, with Section 3 presenting the results of the review. Section 4 describes the set-up of the empirical evaluation, with the results presented in Section 5. Finally, overall conclusions and directions for future work are presented in Section 6.
@@ 2.1 Methodology
We conducted a systematic literature search using the queries “graph analy*” AND “money launder*” and “network analy*” AND “money launder*” on WoS and Scopus, chosen for their high completeness and research credibility [97, 14]. The search is limited to English-language papers published before 2023, excluded recent publications to ensure adequate time for community evaluation.
The selection process included a title and abstract scan, followed by a full-text review. Papers were excluded if they focused on legal aspects or lacked substantial application of network analytics to anti-money laundering (AML). The final set of papers was classified according to seven criteria: publication data, method type, modelling approach, evaluation metrics, main objectives, data type, and network characteristics.
Our review provides both a broad analysis of trends across all selected papers and an in-depth focus on top-cited works (top $10\%$ annually and overall, per Google Scholar). This dual approach highlights general trends and the state-of-the-art (SOTA) in the field. Finally, we examine relevant review papers to contextualize our findings and demonstrate how our work extends existing literature.
Table 1: The (sub-)categories with explanation for the classification of the literature.
@@
Method Type Supervised Application of supervised method
@@
Modelling Method Rule-based Application of specific rules or cut-offs
Manual features Network features having a exact definition
Walk-based Network analysis based on (random) walks
Shallow representation Embeddings typically through matrix factorization or random walks, without employing deep neural networks
Deep representation Embeddings based on deep neural networks, e.g., GNNs
Correlation-based Analysis of correlation of network features with target
Logistic regression Application of logistic regression
Tree-based Application of tree-based methods
SVM-based Application of support vector machines
Neural networks Learning using neural networks
Anomaly detection Application of anomaly or outlier detection
Clustering Features based on community detection or clustering of features
@@
Network Multiple Usage of multiple networks
Inverse/undirected Additional use of network with directions removed or reverted
@@
Figure 1: The search and filtering method with the number of remaining papers between brackets.
@@ 2.2 Results
Figure 1 summarizes the initial search results from WoS and Scopus, including duplicates. After filtering and removing duplicates, 97 papers remained, of which 11 were review papers. Classification based on Table 1 is detailed in Tables LABEL:tab:literature and Table LABEL:tab:literature2, with review papers excluded from category-specific analysis due to their broader scope.
Figure 2 illustrates the increasing trend in AML-related research, with additional category-specific trends shown in Appendix B. Table 2 lists journals and conferences with more than one relevant publication, covering 25 papers; the remaining 72 papers were distributed across other venues with a single publication each.
Figure 2: The number of papers incorporated in the full query and after filtering per year.
Crypto-related research accounts for 26 papers ($26.5\%$ of the total), analysed separately in Appendix C due to its distinct AML setting, data characteristics, and anonymity challenges. These differences are evident in the corresponding plots.
Appendix D presents results for the 20 top-cited papers, summarized in Table 4. The category distribution aligns with the overall literature, with $25\%$ focusing on cryptocurrency, closely matching the full-scope percentage.
Table 2: Journals/conferences that have two or more publications in scope of this review.
@@ 3.1 Review Papers
Ngai et al. [100], Kurshan and Shen [74] and Lokanan [84] focus on financial fraud in general, with less emphasis on anti-money laundering or graph learning. Ngai et al. [100] construct a framework to classify financial fraud, but includes only one paper on money laundering, which was also the only paper on network analytics. Kurshan and Shen [74] give a high-level overview of the general difficulties of machine learning solutions for fighting financial crime. They noted that network-based methods have a lot of potential to process many transactions handled by financial institutions. The paper is closer to a discussion paper than to a systematic literature review. The scope of the review by Lokanan [84] is limited. It deals with both credit card fraud and money laundering, but only incorporates visualisation methods used for outlier detection.
@@
Other review papers deal specifically with AML [75, 115, 42]. The work by Kute et al. [75] is most like ours. The authors give an overview of the methods applied in the literature, but their main topic is interpretability. Hence, their focus is less on the wider AML literature. Semenov et al. [115] discuss the different methods applied to AML. It touches upon different network methods, but it is not a systematic literature review. The framework by Gao and Ye [42] again deals with a wider array of methods, with the authors concluding that network analytics is the most promising. However, the number of papers on network analytics for AML included is limited.
@@
Gerbrands et al. [47] look at the effects of money laundering policies, by measuring the change in network features when a major AML law is introduced. It analyses cluster size and the degree, closeness and betweenness centrality. Hence, the scope is different, since we focus on enhancing anti-money laundering using network features.
@@
A significant number of review papers deal with crypto currency related money laundering [64, 27, 4, 56]. Irwin and Turner [64] describe challenges in detecting money laundering on the blockchain. As a result, the scope is smaller, with a total of 45 unique references. These different challenges are partially addressed by Alarab et al. [4] and Han et al. [56]. The experiment by Alarab et al. [4] compares different supervised machine learning methods, with the statistics of the ego-network of the different nodes in the Bitcoin transaction network [37, 139] as features.
@@
Other papers cover a specific money laundering method and propose tailor-made solutions. Han et al. [56] discuss the application of anomaly detection, based on crime-specific transaction patterns. The paper also covers Ponzi schemes and blackmail campaigns. Their scope is more limited, with a total of 33 references. In the same line, Day [27] gives an overview of different characteristics of financial crime, and uses these to construct knowledge graphs for building legal cases.
We extend on prior work by conducting the first systematic review of network analytics methods for anti-money laundering, including both supervised and unsupervised methods. Our comprehensive scope spans traditional centrality measures to advanced graph neural networks, and addresses money laundering in both fiat and cryptocurrency contexts.
3.2 Full Scope of Papers
We start by analysis the 83 non-review papers. Although network analytics for anti-money laundering has been researched for a long time, especially after 2010 a lot of new research has been published. Figure 2 illustrates this increasing research interest, with an average increase in output of $61\%$ year-on-year for the papers in scope. Below, we dive deeper into the different aspects of these papers.
3.2.1 Methods.
The methods in the literature are supervised, unsupervised, semi-supervised and visualisations. The most popular method is unsupervised learning, followed by supervised learning and visualisations, as shown in Figure B1. The popularity of unsupervised learning may be due to money laundering being very uncommon, resulting in highly skewed data sets with very few labelled observations [3]. This is exacerbated by the limited resources institutions have to investigate the massive amounts of transactions, leaving most transactions not investigated [133, 107]. In addition, criminals try to evade detection by continuously changing tactics [10, 107], making historically suspicious patterns less relevant for future predictions [15]. The performance of (supervised) models trained on historical data can therefore decrease over time [14].
Diving deeper into the different modelling methods presented in Table LABEL:tab:literature2 and Figure B2, we see that half of the papers rely on manual feature engineering. These manual features often involve the classic network centrality measures (degree, betweenness etc.) and summary statistics on the transaction in the nodes’ ego-network. The second is clustering, which is broadly speaking used in two ways. On the one hand, clustering is applied for visualisation to avoid visual cluttering and give the user the possibility to select the level of detail shown. On the other, clusters are used to construct additional features to be used by detection models. Next to manual feature engineering and clustering, many methods can be seen as rule-based. All these insights show that the current literature still heavily relies on expert knowledge (to construct features and/or rules). Hence, there is a large opportunity to further explore and develop more data-driven machine learning methods in the fight against money laundering.
The evolution of methods used over time, as presented in Figure B2(b), shows that the earliest research relied almost exclusively on rule-based systems and manual feature engineering. These remain popular approaches to this day. From 2018 onward, we see the application of deep learning methods based on shallow representations and graph neural networks. Especially deep representations based on GNNs have seen a spike in 2022, the final year covered in this review.
The downstream methods used for classification have also evolved. Tree-based methods, clustering, anomaly detection and logistic regression have been used in earlier literature and are still in use today. However, we see that SVM-based methods were popular until 2019, while neural networks were only used in AML as of 2020. These are often recurrent neural networks to incorporate the temporal aspect of the networks [145, 154, 96, 81].
Key to these methods is also how the authors define the network(s) used. The construction of multiple networks, e.g., same entities but using different relations, to capture richer data is present in 18 papers. Of these papers, five used the undirected network or the one with direction reversed in tandem with the original network. It allows to propagate network information from, e.g., the receiver of a transaction back to the sender.
3.2.2 Objectives.
We identified multiple objectives of the methods for anti-money laundering. Figure B3 shows that the objective is most often detecting suspicious clients, followed by detecting suspicious transactions, money flows and communities. Client classification combines payments made over a longer period and the person’s characteristics. This seems better for detecting money laundering, since it is often done over a longer period, with the aim to make the individual transaction appear normal. Therefore, models trained on a series of transactions, leveraging an individual’s behaviour, can obtain higher performance.
Figure C3 shows that methods for crypto-currency are most often introduced for suspicious transaction detection. Due to the pseudo-anonymity, it is much harder to know which wallets belong to what person [157, 150], making it less feasible to build profiles. Therefore, researchers in AML have also looked at de-anonymisation of the wallets [157, 150, 48]. They use network features to cluster wallets together to see if they belong to the same person or organization.
In recent years, more studies have been done on flow/chain detection, tracking payments over multiple steps/people. Although this requires more computing power, it more closely resembles reality [130, 77]. Hence, methods leveraging these different steps and actors are becoming more widely used.
3.2.3 Evaluation.
The most popular metrics to evaluate the methods are the recall, $F1$-score, precision and time, as shown in Figure B4. We see that training/prediction time was a popular method in earlier studies, but relatively few papers use time in more recent work, which might be a consequence of the increase in computing power. Additionally, the popularity of the precision, recall and $F1$-score is a recent and increasing trend in the literature.
We remark two surprising results in our analysis. First, only one paper used the area under the precision-recall curve (AUC-PR). Due to the high label imbalanced, this metric is shown to be more appropriate than the area under the ROC curve (AUC-ROC) for evaluating performance [26, 103]. Therefore, we would expect it to be more widely used. Second, some papers solely reported accuracy [31, 146, 48, 148]. Given the class imbalance and the dependence on a threshold, accuracy fails in providing an appropriate performance assessment.
3.2.4 Data.
Figure B5 shows that most data sets are open-source, although proprietary data is also used in almost $50\%$ of the papers. Only a minority of papers report their findings using synthetic data. There is, however, a major difference between crypto- and other research. Looking at Figure C1, we see that crypto-research is almost exclusively done via open-source data. This results from the blockchain’s freely accessible distributed ledger [139]. On the other hand, classic transaction data cannot be shared, due to privacy reasons, resulting in half the data sets being proprietary.
Following this last remark, one would expect to have (almost) no research done on open-source data sets for fiat currencies. Figure C1 shows that almost $30\%$ of non-crypto related papers use open-source data. These papers can be categorised in four groups:
  • Methods that are evaluated on non-financial network data, whether or not next to proprietary/synthetic money laundering data: Zhou et al. [156], Ovelgönne et al. [102], Micale et al. [93], Prado-Romero and Gago-Alonso [109]
  • Methods evaluated on government data, assumed or stated to be open-source: Didimo et al. [29], Imanpour et al. [63], Velasco et al. [136], Bahulkar et al. [11], Malm and Bichler [88]
  • Methods constructed to analyse leaked documents/scandals, e.g., the Panama papers: Joaristi et al. [69, 70], Magalingam et al. [85], Adriaens et al. [1], Winiecki et al. [141], Helmy et al. [59], Cheong and Si [23]
  • Other: Nandhini and Das [98], Bhalerao et al. [13], Çağlayan and Bahti̇yar [17]
For the Other category, we have the following. Nandhini and Das [98] say that the data is publicly available, without giving any more information. Bhalerao et al. [13] scraped their data from Hack Forums. Finally, Çağlayan and Bahti̇yar [17] use a Kaggle data set that is based on synthetic data, which we categorise as open-source, since it is made publicly available.
We present the open-source data sets in Table 3. The most widely used are almost all crypto-currency data sets. Raw Bitcoin data was used by ten papers. This raw data is often enhanced using laundering data from other platforms. One such popular platform is WalletExplorer, which is used by three papers. It is less relevant today, since the data has not been maintained since 2016. Another recent data set on Bitcoin, which is becoming more widely used, is the Elliptic data sets [37, 139]. Although only published in 2019, it has already been used in six experiments. After Bitcoin, Ethereum seems to be the second most popular crytpo-currency for AML practices.
The most used data set for fiat currency is the Czech Financial Data set (CFD). This data set does not provide any labels. The papers using this data set introduce additional, synthetic money laundering patterns to test the methods [80, 138]. This is not ideal, as this hinders comparison across research and can introduce detection bias.
As mentioned above, AML methods are also being evaluated on leaked data like the ICIJ Offshore Leaks Database (including, e.g., the Panama-papers) and the Enron data set. However, these are less interesting when evaluating the performance on financial data.
Table 3: The open-source data sets used in the different papers.
Data Set Number References
Raw bitcon 10 [89], [142], [106], [143], [129], [112], [150], [125], [111], [151]
Elliptic Data set 6 [5], [145], [96], [122], [78], [81]
Ethereum XBlock.pro 4 [125], [157], [66], [67]
ICIJ Offshore Leaks Database 4 [70], [44], [69], [141]
WalletExplorer 3 [142], [129], [150]
Czech Financial Data set (CFD) 2 [80], [138]
Enron Data set 2 [85], [1]
RenRenDai 1 [146]
Bloxy 1 [67]
iCOV 1 [47]
Secret Presidential Funds scandal 1 [23]
Tencent QQ 1 [159]
Caviar data set 1 [11]
BitcoinTalk 1 [143]
Ripple Network 1 [19]
2015 Ashley Madison extortion scam 1 [106]
FinCEN 1 [29]
BlockCypher 1 [111]